OpenStudy (anonymous):

Series Solution to Differential equation (x^2+2)y''-xy+y=0 y(0)=-6 y'(0)=4 How would you use the initial conditions ? I found the recurrence formula (i will post below) and I am trying to find the coefficients

OpenStudy (anonymous):

The recurrence formula i found: \[a_{n+2}= \frac{-(n-1)^2a_n}{2(n+2)(n+1)}\]

OpenStudy (anonymous):

I get the following terms ( all odd numbers are =0) \[n=0\hspace{.5cm} a_2=\frac{-a_0}{2*2*1}\] \[n=2 \hspace{0.5cm} a_4=\frac{a_0}{(2*3*4)*4}\] \[n=4 \hspace{0.5cm} a_6=\frac{-9a_0}{(2*3*4*5*6)*8}\] I think the coefficient look something like this but I can't find the relation \[ a_{2n}=\frac{???a_0}{(n+2)! 2^n}\]

OpenStudy (anonymous):

it should be 9 for n=2